A weightlifter is holding a 50.0 lb weight equivalent to 222.4 N,with his forearm.His forearm is positioned at 60°,60° with respect to his upper arm.The forearm is supported by contraction of the biceps muscle which causes a torque around the elbow.Assuming that the tension in biceps acts long the vertical irection given by gravity.What tension must the muscle exert to hold the forearm at the position?What is the force on the elbow joint?Assume that the forearms weight is negligible.Give your answers in SI unit


1
Expert's answer
2021-02-05T11:47:57-0500

Assume that the forearm length is d=0.3m and the bicep is attached b=2cm from the elbow joint. Then the equilibrium of torques would imply



"W\cdot d\text{ cos}30°-T\cdot b=0,\\\space\\
T=W\frac{d\text{ cos}30°}{b}=222.4\frac{0.3\cdot0.86}{0.02}=2889\text{ N}."